OpenStudy (anonymous):

Which values for a and b make the statement (a + 2)(b - 9) = 0 true? a = 2, b = -9 a = -2, b = 9 a = -2, b = -9 a = 2, y = 9

OpenStudy (anonymous):

I love answering your question XD What you have to do is set each parenthesis to 0 a+2 = 0 b-9 = 0 Are you able to solve from here?

OpenStudy (anonymous):

Would it be #1?

OpenStudy (anonymous):

Remember you have to take the -9 and 2 over to the other side

OpenStudy (anonymous):

But when you plug is (9) -9 = 0, the equation is correct When you plug in (-2)+2) = 0, the equation is correct You are left with x = -2 and x = 9

OpenStudy (anonymous):

@KirbyLegs Oh I get it.

OpenStudy (anonymous):

@KirbyLegs Thanks!

OpenStudy (anonymous):

Yay!